# Trust-Region Twisted Policy Improvement

### v2

@@ 2.1 Control as Inference
$p_{\pi}(A_{t}|S_{t},\mathcal{O}_{t:T})$ $=\pi(A_{t}|S_{t})\frac{p(\mathcal{O}_{t:T}|S_{t},A_{t})}{p(\mathcal{O}_{t:T}|S_{t})},$ (2)
$\propto\pi(A_{t}|S_{t})\exp[\ln p(\mathcal{O}_{t:T}|S_{t},A_{t})].$
This connects us to an (expected-reward) maximum-entropy reinforcement learning setting (Toussaint, 2009), since the above exponent admits a soft-Bellman recursion,
$Q^{\pi}_{soft}(S_{t},A_{t})$ $=\ln p(\mathcal{O}_{t:T}|S_{t},A_{t}),$ (3)
$=\ln p(\mathcal{O}_{t}|S_{t},A_{t})+\ln\mathbb{E}\>p(\mathcal{O}_{t+1:T}|S_{t+1}),$
$=R(S_{t},A_{t})+\ln\mathbb{E}\exp V^{\pi}_{soft}(S_{t+1}),$
@@ 2.2 Particle Filter Planning for RL
Intuitively, resampling improves our estimate of Eq. 5 by ‘correcting’ particles with low-likelihood to higher likelihood regions of the target distribution. However, resampling also introduces noise and would ideally not be needed (or to a lesser extent) if our proposal distribution $p_{q}$ produced samples that better match our target. To this end, variational SMC methods (Naesseth et al., 2018) learn a proposal distribution $q_{\theta}$ from sample estimates of the posterior target $\mathbb{E}[\hat{q}^{*}]=q^{*}$. Macfarlane et al. (2024) use this strategy with neural networks to learn $q_{\theta}\approx q^{*}$. Given the neural network iterates $\{\theta_{(n)}\}_{n=1}^{N}$, their method also updates the prior policy $\pi_{(n+1)}\leftarrow q_{\theta_{(n)}}$ at every step $n$. In other words, they combine variational SMC in an EM-loop (Neal & Hinton, 1998) by using the learned proposals $q_{\theta_{(n)}}$ as the prior to regularize the posterior $p_{\pi_{(n+1)}}(A_{t}|S_{t},\mathcal{O}_{t:T})$.
@@ 4 Related Work
Figure 2: Per environment average evaluation curves given a planning budget of $N=K\times m=16$. Shaded regions give 99% two-sided BCa-bootstrap intervals over 30 seeds. The A2C curves do not show intervals as Bonnet et al. (2024) only reported 3 seeds.
@@ 5 Experiments
We performed experiments on the Brax continuous control tasks (Freeman et al., 2021) and Jumanji discrete environments (Bonnet et al., 2024), using the authors’ A2C and SAC results as baselines alongside our PPO implementation (Schulman et al., 2017). Although we compare sample-efficiency to the model-free baselines, this is only for reference since we do not account for the additional observed transitions by the planner. Performance is reported as the average offline return over 128 episodes. Unless otherwise stated, all experiments were repeated (retrained) across 30 seeds, with 99% two-sided BCa-bootstrap confidence intervals (Efron, 1987). Full experimental setup, model training details, and hyperparameters are in Appendix B.
@@ 5.2 Ablations
The main results show that our TRT-SMC method can perform better compared to the baseline model-free, variational SMC, and MCTS methods in terms of sample-efficiency and training runtime. However, we would also like to know: to what extent do each of our separate contributions improve the base method? Therefore, this section quantifies this across the different environments and parameter settings.
@@
We assess the effect of the proposal distribution on the normalized final test performance for the Jumanji environments in the barplot of Figure 4. This compares our TRT-SMC method on a grid of planner depths and particle budgets when using either a twisted proposal distribution with a greediness tolerance of $\alpha=0.1$, or $\alpha=0$ (which only uses the prior $\pi_{\theta}$). Similarly to the main results, at low particle budgets we observe significantly improved performance with the trust-region twisted proposal, but this gap decreases for $\alpha=0$ at higher particle budgets. It also shows that performance scales favorably when the particle budget and the planner depth are in tandem with eachother. This performance improvement at low particle budgets matches our aim of increasing particle-efficiency.
@@ Appendix B Experiment Details
Code and Container package will be released on GitHub only upon publication.

### v3

@@ 2.1 Control as Inference
$p_{\pi}(A_{t}|S_{t},\mathcal{O}_{t:T})\propto\pi(A_{t}|S_{t})\exp[\ln p(\mathcal{O}_{t:T}|S_{t},A_{t})].$ (2)
This connects us to an (expected-reward) maximum-entropy reinforcement learning setting (Toussaint, 2009; Ziebart, 2010), since the exponent admits a soft-Bellman recursion,
$\ln p(\mathcal{O}_{t:T}|S_{t},A_{t})$ $=R(S_{t},A_{t})+\ln\mathbb{E}\>e^{V^{\pi}_{soft}(S_{t+1})},$ (3)
@@ 2.2 Particle Filter Planning for RL
Intuitively, resampling improves our estimate of Eq. 5 by ‘correcting’ particles with low-likelihood to higher likelihood regions of the target distribution. However, resampling also introduces noise and would ideally not be needed (or to a lesser extent) if our proposal distribution $p_{q}$ produced samples that better match our target. To this end, variational SMC methods (Naesseth et al., 2018; Gu et al., 2015) learn a proposal distribution $q_{\theta}$ from sample estimates of the posterior target $\mathbb{E}[\hat{q}^{*}]=q^{*}$. Macfarlane et al. (2024) use this strategy with neural networks to learn $q_{\theta}\approx q^{*}$. Given the neural network iterates $\{\theta_{(n)}\}_{n=1}^{N}$, their method also updates the prior policy $\pi_{(n+1)}\leftarrow q_{\theta_{(n)}}$ at every step $n$. In other words, they combine variational SMC in an EM-loop (Neal & Hinton, 1998) by using the learned proposals $q_{\theta_{(n)}}$ as the prior to regularize the posterior $p_{\pi_{(n+1)}}(A_{t}|S_{t},\mathcal{O}_{t:T})$.
@@ 4 Related Work
Figure 2: Per environment evaluation curves for a planning budget of $N=K\times m=16$. Shaded regions give 99% two-sided BCa-bootstrap intervals over 30 seeds. This plot shows improved sample-efficiency of our method when a highly accurate simulator is available.
@@ 5 Experiments
We performed experiments on the Brax continuous control tasks (Freeman et al., 2021) and Jumanji discrete environments (Bonnet et al., 2024), using the authors’ A2C and SAC results as baselines alongside our PPO implementation (Schulman et al., 2017). Although we compare sample-efficiency to the model-free baselines, this is only for reference since we do not account for the additional observed transitions by the planner in the main results. In other words, we are testing the setting where we assume access to a highly accurate simulator for planning purposes (see Appendix D for additional results that also count the simulator/ planner samples). Performance is reported as the average offline return over 128 episodes. Unless otherwise stated, all experiments were repeated (retrained) across 30 seeds, with 99% two-sided BCa-bootstrap confidence intervals (Efron, 1987). Full experimental setup, model training details, and hyperparameters are in Appendix B.
@@ 5.2 Ablations
The main results show that our TRT-SMC method can perform better compared to the baseline model-free, variational SMC, and MCTS methods in terms of sample-efficiency and training runtime. However, we also want to asses: to what extent do each of our separate contributions improve the base method? Therefore, this section quantifies this across the different environments and parameter settings.
@@
Firstly, we assess the interplay of the proposal distribution with the planning budget through the aggregated final test performance on the Jumanji environments in Figure 4. This compares our TRT-SMC method when using either a twisted proposal distribution with a greediness tolerance of $\alpha=0.1$, or $\alpha=0$ (which only uses the prior $\pi_{\theta}$). Similarly to the main results, at low particle budgets we observe significantly improved performance with the trust-region twisted proposal, but this gap decreases for $\alpha=0$ at higher particle budgets. It also shows that performance scales favorably when the particle budget and the planner depth are in tandem with eachother. This performance improvement at low particle budgets matches our aim of increasing particle-efficiency.
@@
Figure 5: Sensitivity plot for the adaptive trust-region parameter $\alpha$ over increasing planning budgets $N=K\cdot m$ (where $K=m$). The $y$-axis indicates the normalized final test performance, aggregated across all tested environments. The pattern shows that at small planning budgets, a proposal distribution which accounts for the predicted $Q^{\pi}_{\theta}$ performs marginally better.
Secondly, we evaluated our TRT-SMC method varying the $\alpha\in[0,1]$ parameter and uniformly scaling up the budget and depth (as done in Figure 3). We aggregated the normalized final test results over all tested environments as given by the sensitivity plot in Figure 5. Although we show the aggregated results here, we found that this pattern highly depends on the specific environment, we show the individual results in Appendix D. In essence, these results show that mixing between the prior $\pi_{\theta}$ and maximizing the predicted state-action value ${Q}_{\theta}^{\pi}$ improves performance compared to completely relying on either of them.
@@ Appendix B Experiment Details
Our code can be found at https://github.com/joeryjoery/trtpi.
@@ Appendix D Supplementary Results
Figure 8: Main results from Figure 2 with the adjusted $x$-axis to account for additional true environment samples during planning. In terms of sample-efficiency this gives a stark contrast to the result of the main paper, and shows that the model-free methods are more sample efficient. However, methods that utilize planning can always utilize an accurate simulator to skew the $x$-axis similarly to the result of Figure 2.
Figure 9: Sensitivity plot for the adaptive trust-region parameter $\alpha$ split for the discrete (Jumanji) and continuous (brax) environments. This splits up the results shown in Figure 5 from the main paper.